OpenStudy (sonyalee77):

Sales at a certain department store follow the model y=85-13.254x where y is the total sales in thousands of dollars and x is the number of years after 2001. What was the first year that sales fell below $50,000? a. 1999 c. 2004 b. 1998 d. 2005

OpenStudy (sonyalee77):

@austinL @texaschic101

OpenStudy (texaschic101):

y = 85 - 13.254x (I am trying 2004, which is 3 years after 2001) y = 85 - 13.254(3) y = 85 - 39.762 y = 45.238 I am gonna say this one....your question is a little unclear because the equation is set up for the years after 2001....your answer choices contain answers that are before 2001. So I automatically crossed them off as being an answer. Maybe I am wrong.....you might need to get a second opinion, but I am saying answer choice C. 2004